1. Women have head circumferences that are normally distrusted with a mean of 22.65 in and a standard deviation of 0.80 in. A company makes hats with a circumference from 21.5 in to 24.25 inches. What is the percentage of women that will be able to wear their hats? (Round to 2 decimal places)
11:17 PM iPad a mathxl.com Home Navy Federal Credit... Email Profess Dev... Mail-snfulenwider... Math 1530-113 $19 Sier Homework: HW 1.3,1.4,1.5 Score: 0 of 1 pt 1.3.3 1 of 26 (0 complete) What does it mean when sampling is done without replacement? Choose the correct answer below O A. Once an individual is selected, the individual can be selected again. O B. Once a sample is taken, those individuals cannot be selected for any other samples. O C. Once a sample is taken, those individuals are no longer considered part of the population. O D. Once an individual is selected, the individual cannot be selected again
Answer #1

Ans:

1)

Mean=22.65

standard deviation=0.80

z(21.5)=(21.5-22.65)/0.80=-1.4375

z(24.25)=(24.25-22.65)/0.80=2

P(-1.4375<z<2)=P(z<2)-P(z<-1.4375)

=0.9773-0.0753=0.9020 or 90.20%

2)Option D is correct.

Once an individual is selected, the individual cannot be selected again